OpenStudy (anonymous):

verify : sin(theta)+cos(theta) = (sin(theta)/1-cot(theta)) + (cos(theta)/1-tan(theta))

OpenStudy (anonymous):

\[a+b=\frac{b}{1-\frac{a}{b}}+\frac{a}{1-\frac{b}{a}}\] for starters